Jackson Inc. produces leather handbags. The production budget for the next four months is: July 5,700 units, August 7,200 units, September 8,300 units, October 8,200 units. Each handbag requires 0.6 square meters of leather. Jackson Inc.’s leather inventory policy is 25% of next month’s production needs. On July 1 leather inventory was expected to be 855 square meters. Leather is expected to cost $8.00 per square meter in June, but go up to $9.00 per square meter in July. What is the expected cost of leather purchases in July?

  • $31,305

  • $33,805

  • $33,705

  • $32,805
Answer #1

Solution:

July production needs = 5700 units* 0.6 square meter per unit = 3420

July beginning leather inventory = 855 square meters

July ending leather inventory = August product needs *25% = 7200*0.6 * 25% = 1080 sqare meters

July leather purchases = Ending inventory + July production needs - beginning inventory

= 1080 + 3420 - 855 = 3645 square meters

Cost of leather purchases in July = 3645* $9.00 = $32,805

Hence last option is correct.